# Robust Molecular Property Prediction via Densifying Scarce Labeled Data

### v3

@@ 1 Introduction
Molecular property prediction plays a central role in drug discovery pipelines, enabling researchers to prioritize compounds for costly and time-consuming experimental validation. Accurate computational models have the potential to dramatically accelerate early-stage discovery by predicting critical attributes such as bioactivity, toxicity, and solubility before synthesis (Schneider, 2018; Vamathevan et al., 2019). However, building reliable predictive models generalizing to novel, unseen compounds remains a fundamental challenge.
Standard molecular property prediction models tend to rely heavily on patterns observed within the training distribution, resulting in poor generalization to out-of-distribution compounds (Klarner et al., 2023; Ovadia et al., 2019; Koh et al., 2021). In drug discovery, this limitation is particularly problematic, since the compounds most crucial for advancing research often lie far beyond the chemical spaces represented during training (Lee et al., 2023). The resulting covariate shift introduces significant obstacles to reliable prediction, with models frequently producing unstable outputs when extrapolating to new regions of chemical space. Further compounding these challenges, experimental validation of molecular properties is both costly and resource-intensive, leading to a scarcity of labeled data and increasing reliance on computational exploration (Altae-Tran et al., 2017). Also, available labeled data is typically concentrated in narrow regions of chemical space, introducing bias that hampers generalization to unseen compounds (Klarner et al., 2023).
While vast collections of unlabeled molecular structures are readily available (Sterling & Irwin, 2015; Kim et al., 2021), offering rich information about the structure of chemical space, existing methods often fail to fully exploit this resource to improve generalization (Klarner et al., 2023). Therefore, we propose a novel bilevel optimization method that leverages unlabeled data to densify the scarce train dataset and guide the model toward sensible behavior in unexplored regions of chemical space. Our code can be found at https://github.com/JinA0218/drugood-densify.
@@ 2 Methodology
Scarce Data Densification with Unlabeled Data To address this, we propose a bilevel optimization framework that interpolates the training distribution $\mathcal{D}_{\text{train}}$ with an exogenous distribution $\mathcal{D}_{\text{unlabeled}}$. Our objective is to leverage the cheaper and more abundant distribution $\mathcal{D}_{\text{unlabeled}}$ to densify the scarce labeled distribution $\mathcal{D}_{\text{train}}$ in a way that encourages the model to generalize robustly under covariate shift, particularly in out-of-distribution scenarios where we have no label information and therefore high uncertainty. For this, we utilize subsets of $\mathcal{D}_{\text{unlabeled}}$ as $\mathcal{D}_{\text{context}}$ and $\mathcal{D}_{\text{mvalid}}$, where $\mathcal{D}_{\text{context}}$ is a domain-informed external task distribution for interpolating to $\mathcal{D}_{\text{train}}$, and $\mathcal{D}_{\text{mvalid}}$ is a meta-validation set used to guide the interpolation function. Inspired by (Lee et al., 2024), we introduce a permutation invariant learnable set function (Zaheer et al., 2017; Lee et al., 2019) $\mu_{\lambda}$ as a mixer (interpolator), which learns to mix each point from $x_{i}\sim\mathcal{D}_{\text{train}}$ with the context points $\{c_{ij}\}_{j=1}^{m_{i}}$ in a way that densifies $\mathcal{D}_{\text{train}}$, where
@@
We utilize bilevel optimization for training meta-learner $f_{\theta_{l}}$, and treat the set function parameter $\mu_{\lambda}$ as a hyperparameter to be optimized in the outer loop (Lorraine et al., 2019). As shown in Table 2 (w/o bilevel optimization), simply optimizing the meta-learner parameters $\theta$ and the set function parameters $\lambda$ jointly can lead to overfitting to the task distribution and harms test-time generalization. Following the setting of (Lorraine et al., 2019), during training, we only update the parameter $\theta$ in the inner loop and we only update the parameter $\lambda$ in the outer loop (see Figure 9(b) for the detailed model structure of the bilevel optimization).
@@
In the outer loop, we train the set encoder using hypergradient (Lorraine et al., 2019), which aims to minimize a meta-validation loss $L_{V}(\lambda,\theta^{*}(\lambda))$, where the model parameters $\theta^{*}(\lambda)$ are the solution to the inner training objective $\theta^{*}(\lambda)=\arg\min_{\theta}L_{T}(\theta,\lambda)$ and $L_{T}(\theta,\lambda)$ denotes the training loss computed on a labeled dataset $\mathcal{D}_{\text{train}}$, potentially regularized or influenced by the hyperparameters $\lambda$. This inner loss $L_{T}$ governs the optimization of model parameters $\theta$, while $L_{V}$ evaluates generalization performance on a meta-validation set $\mathcal{D}_{\text{mvalid}}$, and guides the update of $\lambda$. In order for the hypergradient to train the set encoder $\mu_{\lambda}$ in a way that guides the overall model $\hat{f}_{\theta,\lambda}$ toward robustness under covariate shift, we construct $\mathcal{D}_{\text{mvalid}}$ as
@@
Luckily, the inverse Hessian vector product can be approximated and computed efficiently using automatic differentiation and Neuman series iterations (Lorraine et al., 2019). In Section 4, we display how the model $\hat{f}_{\theta,\lambda}$ effectively uses $\mathcal{D}_{\text{context}}$ to densify the data distribution around $\mathcal{D}_{\text{train}}$.
@@ 3 Experiments
We evaluate on the Merck Molecular Activity Challenge (Merck) dataset (Kulkarni et al., 2012), a benchmark for molecular property prediction. The dataset contains bioactivity measurements for therapeutic targets, with molecules represented by high-dimensional chemical descriptors. It comprises 15 distinct datasets with a regression target predicting a molecular activity measurement in (possibly) different units, enabling evaluation across multiple distribution shifts. The Merck dataset reflects real-world drug discovery scenarios where promising candidates lie outside known chemical spaces, making generalization challenging. Molecular inputs are encoded either as bit vectors, indicating the presence or absence of substructures, or count vectors, which reflect the frequency of substructures—providing varying granularity of chemical information.
Following the prior work (Klarner et al., 2023), for $\mathcal{D}_{\text{train}}$, we focus on the three subsets (HIVPORT, DPP4, NK1) that exhibit the greatest distributional shift from and treat the remaining 12 datasets as $\mathcal{D}_{\text{unlabeled}}$ and exclude their labels from the training process. For fair comparison with baselines, we randomly select some portion ($K$ for each minibatch) of $\mathcal{D}_{\text{unlabeled}}$ as $\mathcal{D}_{\text{mvalid}}$ and sample $\mathcal{D}_{\text{context}}$ from $\mathcal{D}_{\text{unlabeled}}$ ($m_{i}$ for each minibatch). As $\mathcal{D}_{\text{mvalid}}$ is randomly labeled, we ensure that all baselines see the same amount of labeled data.
In addition to the baselines presented in (Klarner et al., 2023), we further implement and evaluate two related state-of-the-art interpolation methods—Mixup (Zhang et al., 2017) and Manifold Mixup (Verma et al., 2019)—to empirically demonstrate the capability of our set function $\mu_{\lambda}$. As mixup performs the mixing procedure in the input space, there are no parameters to train for bilevel optimization. We therefore report two variants of mixup (plain and outlier exposure) where outlier exposure is exposed to the same $\mathcal{D}_{\text{mvalid}}$ with random labels as described above. Manifold mixup performs the mixup procedure in the feature space between layers, therefore we can straightforwardly apply our bilevel optimization procedure to manifold mixup. Therefore, we report two variants of manifold mixup (plain and bilevel optimization).
@@
For Mixup, we set $l_{\text{mix}}=1$ so we mix $\{x_{i}^{{(l_{\text{1}})}},C_{i}^{{(l_{\text{1}})}}\}$ and for Manifold Mixup, we set $l_{\text{mix}}>1$, so we mix $\{x_{i}^{{(l_{\text{mix}})}},C_{i}^{{(l_{\text{mix}})}}\}$ after passing at least one layer $f_{\theta_{l}}$ (For further implementation details, please see Section A.1). When evaluating with bilevel optimization, we follow our setting and update $\theta$ and $\lambda$ separately in each loop using hypergrad (Lorraine et al., 2019), and for the setting without bilevel optimization, we jointly update $\theta$ and $\lambda$ using a single optimizer. As shown in Table 1, when either leveraging Deepsets (Zaheer et al., 2017) or Set Transformer (Lee et al., 2019) as the mixer $\mu_{\lambda}$, our method significantly improves MSE compared to the baselines, demonstrating greater robustness of our model toward covariate shift.
Furthermore, we present a series of ablation studies in Table 2 to analyze the contributions of each component of our method (Implementation details can be found in Section A.2). Specifically, to validate the effectiveness of $\mathcal{D}_{\text{context}}$ and our bilevel optimization (Lorraine et al., 2019) under $y_{i,k}^{(\text{mvalid})}\sim\mathcal{N}(0,1)$, we compare our approach against the following variants: (1) using $\mathcal{D}_{\text{context}}$ without bilevel optimization, (2) excluding both $\mathcal{D}_{\text{context}}$ and bilevel optimization, and (3) using the real label for $y_{i,k}^{(\text{mvalid})}$. The results indicate that, in order to achieve robust generalization across all datasets, it is crucial to leverage bilevel optimization with $\mathcal{D}_{\text{context}}$. Especially, the dataset HIVPROT, which has the highest degree of label and covariate shift compared to others (Klarner et al., 2023) (Table 3), showed a detrimental performance drop ($62.6\%\text{–}67.3\%$ for the count vector, and $63.8\%\text{–}65.3\%$ for the bit vector) in settings (1,2). Moreover, using $y_{i,k}^{(\text{mvalid})}\sim\mathcal{N}(0,1)$ in $L_{V}$ for training the set function $\mu_{\lambda}$ resulted in comparable or better performance than using the true label associated with $x_{i,k}^{(\text{mvalid})}$, validating our choice of $L_{V}$ and its contribution to improved robustness under covariate shift.
@@ 4 Analysis
To visually assess how our model leverages $\mathcal{D}_{\text{context}}$ to densify $\mathcal{D}_{\text{train}}$, and to compare model behaviors on out-of-distribution (OOD) inputs, we visualize the penultimate-layer embeddings $\mathbf{z}_{i}=x_{i}^{(L-1)}$ , using t-SNE (Van der Maaten & Hinton, 2008) across three models that interpolate $\mathcal{D}_{\text{train}}$ with $\mathcal{D}_{\text{context}}$: Ours, Mixup (w/ outlier exposure), and Manifold Mixup (w/ bilevel optim.), all trained on the DPP4 bit-vector dataset. We randomly sample $x_{i},y_{i}\sim\mathcal{D}_{\text{train}}$, $C_{i}=\{c_{ij}\}_{j=1}^{100}\sim\mathcal{D}_{\text{context}}$, $x_{i}^{\text{ood}}\sim\mathcal{D}_{\text{ood}}$, where $\mathcal{D}_{\text{ood}}$ signifies an out-of-distribution dataset which was never seen during training. In Figure 2, each color indicates a penultimate-layer embedding $\mathbf{z}\in\mathbf{Z}$ under different input configurations. We denote the embeddings as follows: $\mathbf{Z}_{\text{joint}}$ for orange ($\{x_{i},C_{i}\}$), $\mathbf{Z}_{\text{input}}$ for blue ($x_{i}$ alone), $\mathbf{Z}_{\text{context}}$ for green ($C_{i}$ alone), and $\mathbf{Z}_{\text{ood}}$ for purple (OOD samples $x_{i}^{\text{ood}}$). Additional t-SNE results and details are in Appendix B.
@@ 5 Conclusion
Appendix A Experimental Details
@@
A.1 Implementation details of the baselines (Table 1)
In Table 1, we report the performance of $L_{1}$-Regression, $L_{2}$-Regression, Random Forest, MLP, and Q-SAVI by referring to the results from (Klarner et al., 2023). For Mixup (plain and with outlier exposure) and Manifold Mixup (plain and with bilevel optimization), we replace only our set function $\mu_{\lambda}$ with the linear interpolation $\mu_{\alpha}^{\text{lin}}$ and reduction function, defined as
@@ A.1 Implementation details of the baselines (Table 1)
Mixup (w/ outlier exposure). To compare Mixup with outlier exposure, we implemented a variant of Mixup which is trained with $(x_{i},y_{i})\sim\mathcal{D}_{\text{train}}$, $\{c_{ij}\}_{j=1}^{m_{i}}\sim\mathcal{D}_{\text{context}}$ and $\{x_{i,k}^{(\text{mvalid})}\}_{k=1}^{K}\sim\mathcal{D}_{\text{unlabeled}},\;\;y_{i,k}^{(\text{mvalid})}\sim\mathcal{N}(0,1)$. As explained in Section 3, since it peforms the mixing proceduce in the input space there are no parameters to train, it is impossible to apply bilevel optimization. Therefore, when computing the loss, we do not follow the operation held in hypergradient (Lorraine et al., 2019) but instead add train loss and $L_{V}$, and update the model parameter with a single optimizer.
@@
A.2 Implementation details of the ablation study (Table 2)

### v4

@@ 1 Introduction
Molecular property prediction plays a central role in drug discovery pipelines, enabling researchers to prioritize compounds for costly and time-consuming experimental validation. Accurate computational models have the potential to dramatically accelerate early-stage discovery by predicting critical attributes such as bioactivity, toxicity, and solubility before synthesis (schneider2018automating; vamathevan2019applications). However, building reliable predictive models generalizing to novel, unseen compounds remains a fundamental challenge.
Standard molecular property prediction models tend to rely heavily on patterns observed within the training distribution, resulting in poor generalization to out-of-distribution compounds (qsavi; ovadia; koh). In drug discovery, this limitation is particularly problematic, since the compounds most crucial for advancing research often lie far beyond the chemical spaces represented during training (mood). The resulting covariate shift introduces significant obstacles to reliable prediction, with models frequently producing unstable outputs when extrapolating to new regions of chemical space. Further compounding these challenges, experimental validation of molecular properties is both costly and resource-intensive, leading to a scarcity of labeled data and increasing reliance on computational exploration (altae2017low). Also, available labeled data is typically concentrated in narrow regions of chemical space, introducing bias that hampers generalization to unseen compounds (qsavi).
While vast collections of unlabeled molecular structures are readily available (zinc; pubchem), offering rich information about the structure of chemical space, existing methods often fail to fully exploit this resource to improve generalization (qsavi). Therefore, we propose a novel bilevel optimization method that leverages unlabeled data to densify the scarce train dataset and guide the model toward sensible behavior in unexplored regions of chemical space. Our code can be found at https://github.com/JinA0218/drugood-densify.
@@ 2 Methodology
Scarce Data Densification with Unlabeled Data To address this, we propose a bilevel optimization framework that interpolates the training distribution $\mathcal{D}_{\text{train}}$ with an exogenous distribution $\mathcal{D}_{\text{unlabeled}}$. Our objective is to leverage the cheaper and more abundant distribution $\mathcal{D}_{\text{unlabeled}}$ to densify the scarce labeled distribution $\mathcal{D}_{\text{train}}$ in a way that encourages the model to generalize robustly under covariate shift, particularly in out-of-distribution scenarios where we have no label information and therefore high uncertainty. For this, we utilize subsets of $\mathcal{D}_{\text{unlabeled}}$ as $\mathcal{D}_{\text{context}}$ and $\mathcal{D}_{\text{mvalid}}$, where $\mathcal{D}_{\text{context}}$ is a domain-informed external task distribution for interpolating to $\mathcal{D}_{\text{train}}$, and $\mathcal{D}_{\text{mvalid}}$ is a meta-validation set used to guide the interpolation function. Inspired by (metainterpolation), we introduce a permutation invariant learnable set function (deepsets; set-transformer) $\mu_{\lambda}$ as a mixer (interpolator), which learns to mix each point from $x_{i}\sim\mathcal{D}_{\text{train}}$ with the context points $\{c_{ij}\}_{j=1}^{m_{i}}$ in a way that densifies $\mathcal{D}_{\text{train}}$, where
@@
We utilize bilevel optimization for training meta-learner $f_{\theta_{l}}$, and treat the set function parameter $\mu_{\lambda}$ as a hyperparameter to be optimized in the outer loop (hypergrad). As shown in Table 2 (w/o bilevel optimization), simply optimizing the meta-learner parameters $\theta$ and the set function parameters $\lambda$ jointly can lead to overfitting to the task distribution and harms test-time generalization. Following the setting of (hypergrad), during training, we only update the parameter $\theta$ in the inner loop and we only update the parameter $\lambda$ in the outer loop (see Figure 3(b) for the detailed model structure of the bilevel optimization).
@@
In the outer loop, we train the set encoder using hypergradient (hypergrad), which aims to minimize a meta-validation loss $L_{V}(\lambda,\theta^{*}(\lambda))$, where the model parameters $\theta^{*}(\lambda)$ are the solution to the inner training objective $\theta^{*}(\lambda)=\arg\min_{\theta}L_{T}(\theta,\lambda)$ and $L_{T}(\theta,\lambda)$ denotes the training loss computed on a labeled dataset $\mathcal{D}_{\text{train}}$, potentially regularized or influenced by the hyperparameters $\lambda$. This inner loss $L_{T}$ governs the optimization of model parameters $\theta$, while $L_{V}$ evaluates generalization performance on a meta-validation set $\mathcal{D}_{\text{mvalid}}$, and guides the update of $\lambda$. In order for the hypergradient to train the set encoder $\mu_{\lambda}$ in a way that guides the overall model $\hat{f}_{\theta,\lambda}$ toward robustness under covariate shift, we construct $\mathcal{D}_{\text{mvalid}}$ as
@@
Luckily, the inverse Hessian vector product can be approximated and computed efficiently using automatic differentiation and Neuman series iterations (hypergrad). In Section 4, we display how the model $\hat{f}_{\theta,\lambda}$ effectively uses $\mathcal{D}_{\text{context}}$ to densify the data distribution around $\mathcal{D}_{\text{train}}$.
@@ 3 Experiments
We evaluate on the Merck Molecular Activity Challenge (Merck) dataset (merck_activity), a benchmark for molecular property prediction. The dataset contains bioactivity measurements for therapeutic targets, with molecules represented by high-dimensional chemical descriptors. It comprises 15 distinct datasets with a regression target predicting a molecular activity measurement in (possibly) different units, enabling evaluation across multiple distribution shifts. The Merck dataset reflects real-world drug discovery scenarios where promising candidates lie outside known chemical spaces, making generalization challenging. Molecular inputs are encoded either as bit vectors, indicating the presence or absence of substructures, or count vectors, which reflect the frequency of substructures—providing varying granularity of chemical information.
Following the prior work (qsavi), for $\mathcal{D}_{\text{train}}$, we focus on the three subsets (HIVPORT, DPP4, NK1) that exhibit the greatest distributional shift from and treat the remaining 12 datasets as $\mathcal{D}_{\text{unlabeled}}$ and exclude their labels from the training process. For fair comparison with baselines, we randomly select some portion ($K$ for each minibatch) of $\mathcal{D}_{\text{unlabeled}}$ as $\mathcal{D}_{\text{mvalid}}$ and sample $\mathcal{D}_{\text{context}}$ from $\mathcal{D}_{\text{unlabeled}}$ ($m_{i}$ for each minibatch). As $\mathcal{D}_{\text{mvalid}}$ is randomly labeled, we ensure that all baselines see the same amount of labeled data.
In addition to the baselines presented in (qsavi), we further implement and evaluate two related state-of-the-art interpolation methods—Mixup (mixup) and Manifold Mixup (manifold_mixup)—to empirically demonstrate the capability of our set function $\mu_{\lambda}$. As mixup performs the mixing procedure in the input space, there are no parameters to train for bilevel optimization. We therefore report two variants of mixup (plain and outlier exposure) where outlier exposure is exposed to the same $\mathcal{D}_{\text{mvalid}}$ with random labels as described above. Manifold mixup performs the mixup procedure in the feature space between layers, therefore we can straightforwardly apply our bilevel optimization procedure to manifold mixup. Therefore, we report two variants of manifold mixup (plain and bilevel optimization).
@@
For Mixup, we set $l_{\text{mix}}=1$ so we mix $\{x_{i}^{{(l_{\text{1}})}},C_{i}^{{(l_{\text{1}})}}\}$ and for Manifold Mixup, we set $l_{\text{mix}}>1$, so we mix $\{x_{i}^{{(l_{\text{mix}})}},C_{i}^{{(l_{\text{mix}})}}\}$ after passing at least one layer $f_{\theta_{l}}$ (For further implementation details, please see Section B.1). When evaluating with bilevel optimization, we follow our setting and update $\theta$ and $\lambda$ separately in each loop using hypergrad (hypergrad), and for the setting without bilevel optimization, we jointly update $\theta$ and $\lambda$ using a single optimizer. As shown in Table 1, when either leveraging Deepsets (deepsets) or Set Transformer (set-transformer) as the mixer $\mu_{\lambda}$, our method significantly improves MSE compared to the baselines, demonstrating greater robustness of our model toward covariate shift.
Furthermore, we present a series of ablation studies in Table 2 to analyze the contributions of each component of our method (Implementation details can be found in Section B.2). Specifically, to validate the effectiveness of $\mathcal{D}_{\text{context}}$ and our bilevel optimization (hypergrad) under $y_{i,k}^{(\text{mvalid})}\sim\mathcal{N}(0,1)$, we compare our approach against the following variants: (1) using $\mathcal{D}_{\text{context}}$ without bilevel optimization, (2) excluding both $\mathcal{D}_{\text{context}}$ and bilevel optimization, and (3) using the real label for $y_{i,k}^{(\text{mvalid})}$. The results indicate that, in order to achieve robust generalization across all datasets, it is crucial to leverage bilevel optimization with $\mathcal{D}_{\text{context}}$. Especially, the dataset HIVPROT, which has the highest degree of label and covariate shift compared to others (qsavi) (Table 3), showed a detrimental performance drop ($62.6\%\text{–}67.3\%$ for the count vector, and $63.8\%\text{–}65.3\%$ for the bit vector) in settings (1,2). Moreover, using $y_{i,k}^{(\text{mvalid})}\sim\mathcal{N}(0,1)$ in $L_{V}$ for training the set function $\mu_{\lambda}$ resulted in comparable or better performance than using the true label associated with $x_{i,k}^{(\text{mvalid})}$, validating our choice of $L_{V}$ and its contribution to improved robustness under covariate shift.
@@ 4 Analysis
To visually assess how our model leverages $\mathcal{D}_{\text{context}}$ to densify $\mathcal{D}_{\text{train}}$, and to compare model behaviors on out-of-distribution (OOD) inputs, we visualize the penultimate-layer embeddings $\mathbf{z}_{i}=x_{i}^{(L-1)}$ , using t-SNE (tsne) across three models that interpolate $\mathcal{D}_{\text{train}}$ with $\mathcal{D}_{\text{context}}$: Ours, Mixup (w/ outlier exposure), and Manifold Mixup (w/ bilevel optim.), all trained on the DPP4 bit-vector dataset. We randomly sample $x_{i},y_{i}\sim\mathcal{D}_{\text{train}}$, $C_{i}=\{c_{ij}\}_{j=1}^{100}\sim\mathcal{D}_{\text{context}}$, $x_{i}^{\text{ood}}\sim\mathcal{D}_{\text{ood}}$, where $\mathcal{D}_{\text{ood}}$ signifies an out-of-distribution dataset which was never seen during training. In Figure 2, each color indicates a penultimate-layer embedding $\mathbf{z}\in\mathbf{Z}$ under different input configurations. We denote the embeddings as follows: $\mathbf{Z}_{\text{joint}}$ for orange ($\{x_{i},C_{i}\}$), $\mathbf{Z}_{\text{input}}$ for blue ($x_{i}$ alone), $\mathbf{Z}_{\text{context}}$ for green ($C_{i}$ alone), and $\mathbf{Z}_{\text{ood}}$ for purple (OOD samples $x_{i}^{\text{ood}}$). Additional t-SNE results and details are in Appendix C.
@@ 5 Conclusion
Appendix A Model Structure Overview
(a) Training and Test-time Framework
(b) Bilevel Optim. during Training
Appendix B Experimental Details
@@
B.1 Implementation details of the baselines (Table 1)
In Table 1, we report the performance of $L_{1}$-Regression, $L_{2}$-Regression, Random Forest, MLP, and Q-SAVI by referring to the results from (qsavi). For Mixup (plain and with outlier exposure) and Manifold Mixup (plain and with bilevel optimization), we replace only our set function $\mu_{\lambda}$ with the linear interpolation $\mu_{\alpha}^{\text{lin}}$ and reduction function, defined as
@@ B.1 Implementation details of the baselines (Table 1)
Mixup (w/ outlier exposure). To compare Mixup with outlier exposure, we implemented a variant of Mixup which is trained with $(x_{i},y_{i})\sim\mathcal{D}_{\text{train}}$, $\{c_{ij}\}_{j=1}^{m_{i}}\sim\mathcal{D}_{\text{context}}$ and $\{x_{i,k}^{(\text{mvalid})}\}_{k=1}^{K}\sim\mathcal{D}_{\text{unlabeled}},\;\;y_{i,k}^{(\text{mvalid})}\sim\mathcal{N}(0,1)$. As explained in Section 3, since it peforms the mixing proceduce in the input space there are no parameters to train, it is impossible to apply bilevel optimization. Therefore, when computing the loss, we do not follow the operation held in hypergradient (hypergrad) but instead add train loss and $L_{V}$, and update the model parameter with a single optimizer.
@@
B.2 Implementation details of the ablation study (Table 2)